OpenStudy (anonymous):

Discuss the differences between the measurement of the surface area and the measurement of the volume of a right triangular prism. Include the following topics in your discussion: • Description of regions that are measured • Units for each type of measure and why the units are different • Examples of when you would want to measure for each type of measure

ganeshie8 (ganeshie8):

some data u can use :- 1. surface area = P*h + 2B P = perimeter of triangular base h = height of prism B = area of triangular base 2. volume = B * h B = area of triangular base h = height of prism

ganeshie8 (ganeshie8):

• Description of regions that are measured 1. for surface area you would measure area of base and multiply it with 2(2B), since you have two triangles one up and the other down. then to find the area of sides, you take perimeter of base and multiply it wid the height (P*h) 2. for volume you would measure the area of base and simply multiply it with height (B*h)

ganeshie8 (ganeshie8):

• Units for each type of measure and why the units are different i want you work/try this out.

OpenStudy (anonymous):

Thanks I didn't know you were here

ganeshie8 (ganeshie8):

:) il be here for a while see if u can figure out rest of this activity. i can proof read....

OpenStudy (anonymous):

Hey I need help on this one also. Can you help me out on the unit part?